OpenStudy (anonymous):

Karthik spent 35 dollars of his allowence and 5 dollars on charity. If the number of dollars he spent is porportional to the number of dollars he gives to charity, how much of a 100 dollar allowence will he give to charity?

OpenStudy (anonymous):

amount he spent: $40 ratio spent to amount given to charity: 8:1 you can equate fractions 40/5 = 100/x cross multiply 40x = 500 x = 12.5